UrGirlLily:

Solve for z. Reduce any fractions to lowest terms. Don't round your answer, and don't use mixed fractions. -4z+31≥ 17z + 23

UrGirlLily:

Pleaseee helpp!

rocfp:

I dont even know..

UrGirlLily:

bruh then please dont answer, im really struggling

rocfp:

oop

MrMudd183:

What do you think the first step to do is?

MrMudd183:

The first thing I would do is combine the like terms from each side of the equation. For example, if there is -3x+5>9x+2 You would combine -3x and 9x together, as well as combine 5 and 2 Then the equation would become 3>12x After that, all you need to do is divide 12 on both sides to get x by itself

supie:

I politely disagree with the previous reply We have \(−4z+31≥17z+23\) And the first step would be to get `17z` and subtract it from both sides of the equation which would give us \(−21z+31≥23\) Basically what we just did was \(−4z+31−17z≥17z+23−17z\) So now we have to subtract 31 from both sides of the euqation which would give us \(−21z≥−8\) because \(−21z+31−31≥23−31\longrightarrow −21z≥−8\) Then we divide both sides by `-21` \(\LARGE\frac{-21z}{-21}=\frac{-8}{-21}\) So -21z/-21 cancels out so all you have left is -8/-21 so once that is simplified you should get your answer \(\frac{-8}{2-1}=?\) If we did what MrMudd suggested we would be doing \(3>12x\) \(\frac{3}{12}=\frac{12x}{12}\)

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie I politely disagree with the previous reply We have \(−4z+31≥17z+23\) And the first step would be to get `17z` and subtract it from both sides of the equation which would give us \(−21z+31≥23\) Basically what we just did was \(−4z+31−17z≥17z+23−17z\) So now we have to subtract 31 from both sides of the euqation which would give us \(−21z≥−8\) because \(−21z+31−31≥23−31\longrightarrow −21z≥−8\) Then we divide both sides by `-21` \(\LARGE\frac{-21z}{-21}=\frac{-8}{-21}\) So -21z/-21 cancels out so all you have left is -8/-21 so once that is simplified you should get your answer \(\frac{-8}{-21}=?\) If we did what MrMudd suggested we would be doing \(3>12x\) \(\frac{3}{12}=\frac{12x}{12}\) \(\color{#0cbb34}{\text{End of Quote}}\) Fixing a typo.

supie:

@MrMudd183??

MrMudd183:

yeah? sorry....i'm doing my own hw and not fully paying attention. what supie said is more in depth, i was doing a simple explination to try and not cause confusion

supie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MrMudd183 yeah? sorry....i'm doing my own hw and not fully paying attention. what supie said is more in depth, i was doing a simple explination to try and not cause confusion \(\color{#0cbb34}{\text{End of Quote}}\) Yes thanks. Next time if you cant put your full attention on helping someone just don't do it because it seems to be that this time you gave a incorrect answer.

MrMudd183:

I tried to explain in a simple way but yet make it descriptive by doing an example problem

supie:

Oh ok.

UrGirlLily:

ok ok so im confused whats th answer

UrGirlLily:

nvmm